# Learning Task Representations from In-Context Learning

### v1

@@ 5 Experiments
We employ decoder-only autoregressive transformers: GPT-2 [38] for regression tasks and GPT-J [56] for tasks in the language domain, using the default configurations described in [8] and [48], respectively. GPT-2 is configured with 9.5M parameters (256 hidden dimension) across 12 layers and 8 attention heads per layer, while GPT-J features 6B parameters, 28 layers, and 16 attention heads per layer. The GPT-2 models pre-trained on functional regression data follow the training procedure outlined in [8].
@@
For regression data, we evaluate performance on sequences longer than the maximum length seen during pre-training but within the model’s capacity: $T_{\text{train}}<T_{\textbf{v}}\leq T_{\text{max}}$, as described in Section 4.1. We leverage the understanding that positional information is critical for reasoning tasks; when positional encodings for longer sequences are untrained, the model’s reasoning ability—and consequently, its ICL capability—is effectively blocked. This setup offers a controlled environment to evaluate the fidelity of task encoding methods. Importantly, our goal is not to address the length generalization problem but to ensure that task behavior is preserved in the absence of ICL.
In the language domain, we adopt the setup of shuffling the labels of in-context examples to diminish ICL [48]. Notably, the positional encoding approach is unsuitable for the language tasks we consider, as these are inherently word analogy tasks rather than reasoning tasks, to which the theorem in [61] applies. Additionally, following the setup in [48] ensures fair comparisons with [48, 11, 25].
@@
For regression data, as illustrated in Figure 1, we evaluate the models on classes of linear functions, sparse linear functions, and 2-layer ReLU networks [8]. While decision trees have also been explored as a function class, they exhibit less sensitivity to positional encodings due to their feature-based reasoning, which relies on discrete hierarchical rules rather than point-by-point mappings. By grouping multiple inputs to the same label, decision trees largely bypass positional dependencies and are minimally affected by untrained positional encodings. This supports both our argument and the findings [61] that positional information directly affects ICL performance.
For NLP tasks, we focus on single-token generation, as considered in previous work [11, 48]. The tasks are derived from widely studied natural language corpora, including SST-2 [43] and CoNLL-2003 [47]. Further details are provided in Appendix C.3.1.
@@
For functional regression, we include FV [48] and ICV [29] as baselines for comparison. The other methods discussed in Section A were primarily designed for language tasks and are not applicable to regression tasks (see Appendix C.2.6 for details). Although FV and ICV were originally proposed for linguistic applications, we adapted them for regression tasks with several optimizations. Additionally, we consider Low-Rank Optimization (LoRA) [16] for further discussion. For language tasks, we also include Task Vector (TV) [11] and State Vector (SV) with inner optimization [25] alongside FV and ICV.
@@
Details of our experimental setup and additional results are provided in Appendices C and D, respectively. For reproducibility, the code is available in our GitHub repository:
https://github.com/baturaysaglam/ICL-task-repr.
@@ 5.1 Evaluation on Regression Tasks
The loss curves for ICL inference are shown in Figure 3, with additional evaluations under two types of distribution shifts provided in Appendix D.2. While FV offers some benefits in linear regression, ICV negatively impacts the vanilla model’s performance. Despite extensive optimizations, these results indicate that task encoding formulations, originally crafted for language, do not generalize well to other modalities.
Transformer Transformer + LoRA
Transformer + FV (optimized) Transformer + ICV (tuned)
Transformer + LTV (ours)
(a) $T_{\mathbf{v}}=42$
(a) $T_{\mathbf{v}}=56$
(a) $T_{\mathbf{v}}=71$
(a) $T_{\mathbf{v}}=86$
(a) Linear functions ($T_{\text{train}}=41$)
(a) $T_{\mathbf{v}}=42$
(a) $T_{\mathbf{v}}=56$
(a) $T_{\mathbf{v}}=71$
(a) $T_{\mathbf{v}}=86$
(b) Sparse linear functions ($T_{\text{train}}=41$)
(a) $T_{\mathbf{v}}=102$
(a) $T_{\mathbf{v}}=126$
(a) $T_{\mathbf{v}}=151$
(a) $T_{\mathbf{v}}=176$
(c) 2-layer ReLU NN ($T_{\text{train}}=101$)
Figure 3: Squared error on the query input as a function of the number of demonstrations in prompts. Results are averaged over a batch of 256 tasks randomly selected from the same function class. The shaded area represents the 95% confidence interval over the sampled prompts. The dashed line indicates the number of examples the transformer was trained with, while $T_{\mathbf{v}}$ denotes the prompt length used in LTV training.
LTV generalizes to the OOD case while preserving task information.
LTV shows minimal performance difference when trained with prompt lengths near $T_{\text{train}}$. As $T_{\mathbf{v}}$ increases, LTV’s impact grows, achieving optimal performance comparable to the vanilla model at $T=T_{\text{train}}$. While this might raise concerns about the validity of the results, as LTV is specifically designed to handle out-of-distribution (OOD) cases ($T>T_{\text{train}}$), training LTV with prompt lengths slightly exceeding $T_{\text{train}}$—specifically, $T_{\mathbf{v}}=1.37\times T_{\text{train}}$ for linear functions and $T_{\mathbf{v}}=1.25\times T_{\text{train}}$ for 2-layer ReLU networks—proves sufficient to maintain performance. Extending $T_{\mathbf{v}}$ further offers no significant improvements, indicating that only a small portion of the OOD case is needed for LTV to generalize. This suggests that the attention head weights naturally converge to their optimal values with sufficient training data.
@@
The effects of LTV in regression tasks may resemble those of parameter-efficient fine-tuning (PEFT) methods, as only a small number of parameters ($|\Phi|=L\times J$) are trained to adapt the model to previously unlearned scenarios. To evaluate this, we compare LTV with LoRA (rank 8) [16]. While LoRA fine-tunes approximately 196K of the 9.5M GPT-2 parameters, LTV optimizes only 96. Despite LoRA fine-tuning a significantly larger number of parameters on sequences with a maximum length $T_{\text{max}}$, no notable performance improvement is observed. Since LoRA depends on the pre-trained model’s existing positional representations, it cannot address untrained positional encodings. This observation indicates that the task fidelity achieved by LTV is not a result of sole fine-tuning but rather of effective task encoding that can steer the behavior of the model.
@@
No regularization, optimization techniques, dropout, or specific weight initialization were applied to learn the LTV weights. The weights are unbounded and sampled from the standard Gaussian distribution. Visualizing the learned weights as heatmaps revealed they consistently lie within the interval [-3, 3] across tasks. However, no meaningful patterns were observed, suggesting that any underlying structure may be subtle and challenging to detect visually.
Table 1: Accuracy scores (%) for zero-shot and few-shot (5-shot) predictions, averaged across 256 random seeds. $\pm$ represents the margin of error at a 95% confidence level. The highest accuracy is marked in boldface, and the statistically best-performing method is highlighted.
Dataset Prompt Transformer + FV + ICV + SV + TV + LTV (ours)
Abstractive AG News [67] Zero-shot 0.0 $\pm$ 0.0 30.1 $\pm$ 5.7 33.6 $\pm$ 5.6 31.7 $\pm$ 5.7 34.5 $\pm$ 5.0 46.9 $\pm$ 6.2
Antonym [34] Zero-shot 2.3 $\pm$ 1.9 35.2 $\pm$ 5.9 46.4 $\pm$ 5.6 35.2 $\pm$ 6.1 29.6 $\pm$ 6.0 53.5 $\pm$ 6.2
CommonsenseQA [46] Zero-shot 10.2 $\pm$ 3.7 23.0 $\pm$ 5.2 28.9 $\pm$ 5.2 21.7 $\pm$ 1.1 22.0 $\pm$ 4.0 41.0 $\pm$ 6.1
English-French [24] Zero-shot 0.0 $\pm$ 0.0 23.4 $\pm$ 5.2 40.2 $\pm$ 5.3 24.4 $\pm$ 2.7 25.4 $\pm$ 5.1 53.5 $\pm$ 6.2
English-Spanish [24] Zero-shot 0.0 $\pm$ 0.0 9.0 $\pm$ 3.5 16.5 $\pm$ 4.3 9.7 $\pm$ 3.5 9.4 $\pm$ 2.9 20.7 $\pm$ 5.0
Landmark-Country [12] Zero-shot 0.0 $\pm$ 0.0 60.9 $\pm$ 6.0 62.7 $\pm$ 5.1 62.9 $\pm$ 5.6 69.9 $\pm$ 6.0 71.1 $\pm$ 5.6
Person-Instrument [12] Zero-shot 0.0 $\pm$ 0.0 11.3 $\pm$ 3.9 17.1 $\pm$ 4.9 13.1 $\pm$ 2.2 30.1 $\pm$ 3.5 33.2 $\pm$ 5.8
Person-Occupation [12] Zero-shot 0.0 $\pm$ 0.0 4.3 $\pm$ 2.5 17.2 $\pm$ 3.9 9.0 $\pm$ 2.9 8.1 $\pm$ 5.8 36.3 $\pm$ 5.9
Person-Sport [12] Zero-shot 0.0 $\pm$ 0.0 2.7 $\pm$ 2.0 12.9 $\pm$ 3.5 12.5 $\pm$ 3.1 12.6 $\pm$ 3.0 16.0 $\pm$ 4.5
Product-Company [12] Zero-shot 0.0 $\pm$ 0.0 29.3 $\pm$ 5.6 43.0 $\pm$ 5.3 28.2 $\pm$ 4.9 28.7 $\pm$ 5.6 66.4 $\pm$ 5.8
Sentiment Analysis [44] Zero-shot 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 4.8 $\pm$ 1.9 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 10.9 $\pm$ 3.8
Synonym [34] Zero-shot 1.2 $\pm$ 1.3 2.7 $\pm$ 2.0 14.2 $\pm$ 3.6 2.6 $\pm$ 4.6 2.5 $\pm$ 3.3 18.8 $\pm$ 4.8
Extractive NER-person [47] Zero-shot 5.5 $\pm$ 2.8 48.8 $\pm$ 6.2 49.9 $\pm$ 5.8 46.3 $\pm$ 1.3 47.8 $\pm$ 5.3 58.2 $\pm$ 6.1
NER-location [47] Zero-shot 6.6 $\pm$ 3.1 37.9 $\pm$ 6.0 26.4 $\pm$ 5.1 34.6 $\pm$ 3.9 27.2 $\pm$ 4.7 27.3 $\pm$ 5.5
NER-organization [47] Zero-shot 21.9 $\pm$ 5.1 48.8 $\pm$ 6.2 52.5 $\pm$ 6.0 46.1 $\pm$ 1.2 46.6 $\pm$ 4.3 64.1 $\pm$ 5.9
@@ 5.2 Evaluation on Language Tasks
The accuracy scores are reported in Table 1. While “filtered” accuracies was primarily considered in [48], which take into account only the test queries where at least one model responds correctly, we present unfiltered accuracies as a fairer metric, counting all samples regardless of model performance.
@@
In 29 out of 30 evaluations (15 tasks, zero- and few-shot prompts), LTV achieves the highest accuracy and is among the statistically best-performing methods. Furthermore, in 19 out of 30 evaluations, LTV’s performance is statistically superior to all other techniques. We also observe a notable performance by ICV, particularly when the vanilla transformer performs poorly in the zero-shot setting. This effectiveness is attributed to ICV’s approach of crafting unique task vectors for each layer, similar to our method. Overall, these findings confirm that LTV excels in both regression and challenging linguistic tasks, supporting our hypothesis about how ICL tasks are formulated within the transformer architecture.
@@ 5.3 Ablation Studies
The results are reported in Table 2. We also include the results for FV, as it is the only baseline method that does not degrade the vanilla transformer. A detailed description of our experimental methods, along with a figure illustrating the pipeline, is provided in Appendix C.6. Additionally, probability densities of the last hidden state distributions, visualized through histograms, are included in Appendix D.3.
@@
We observe that the divergence between the last hidden state distributions of the LTV-integrated model and the optimal-performing model is generally low and significantly decreases when the LTV training length matches or exceeds the middle length (56 for linear models and 126 for 2-layer NN functions). This suggests that LTV’s effectiveness primarily stems from its ability to dominate the last hidden state and align it with that of the optimal-performing model, thereby facilitating task steering. This observation further implies that task steering may generally be achieved through this mechanism, not exclusively by our method. However, a deeper analysis is required to fully understand this process, which is beyond the scope of this study.
Transformer 1.098 1.110 0.103
+ Function Vector 0.646 0.572 0.223
+ LTV ($T_{\mathbf{v}}=\{41,41,101\}$) 1.000 1.056 0.095
+ LTV ($T_{\mathbf{v}}=\{42,42,102\}$) 0.952 0.984 0.037
+ LTV ($T_{\mathbf{v}}=\{56,56,126\}$) 0.532 0.260 0.016
+ LTV ($T_{\mathbf{v}}=\{71,71,151\}$) 0.460 0.203 0.017
+ LTV ($T_{\mathbf{v}}=\{86,86,176\}$) 0.344 0.199 0.013
+ LTV ($T_{\mathbf{v}}=\{101,101,201\}$) 0.300 0.096 0.038

### v2

@@ 5 Experiments
We employ decoder-only autoregressive transformers: GPT-2 (Radford et al., 2019) for regression tasks and GPT-J (Wang and Komatsuzaki, 2021) for tasks in the language domain, using the default configurations described by Garg et al. (2023) and Todd et al. (2024), respectively. GPT-2 is configured with 9.5M parameters (256 hidden dimension) across 12 layers and 8 attention heads per layer, while GPT-J features 6B parameters, 28 layers, and 16 attention heads per layer. The GPT-2 models pre-trained on functional regression data follow the training procedure outlined by Garg et al. (2023).
@@
For the regression data, we evaluate performance on sequences longer than the maximum length seen during pre-training but within the model’s capacity: $T_{\text{train}}<T_{\textbf{v}}\leq T_{\text{max}}$, as described in Section 4.1. This setup provides a controlled environment to assess the functional restoration of task behavior through causal interventions. In the language domain, we adopt the approach of shuffling labels of in-context examples to diminish ICL (Todd et al., 2024), ensuring fair comparisons with Todd et al. (2024); Hendel et al. (2023); Li et al. (2024).
Transformer Trans. +LoRA Trans. +FV (optimized) Trans. +ICV (tuned) Trans. +LTV
(a) $T_{\mathbf{v}}=56$
(a) $T_{\mathbf{v}}=71$
(a) Linear functions ($T_{\text{train}}=41$)
(a) $T_{\mathbf{v}}=56$
(a) $T_{\mathbf{v}}=71$
(b) Sparse linear functions ($T_{\text{train}}\hskip-2.00749pt=\hskip-1.00374pt41$)
(a) $T_{\mathbf{v}}=126$
(a) $T_{\mathbf{v}}=151$
(c) 2-layer ReLU NN ($T_{\text{train}}=101$)
Figure 3: Squared error on the query input, averaged over a batch of 256 tasks randomly selected from the same function class. The shaded area represents the 95% confidence interval over the sampled prompts. The dashed line indicates the number of examples the transformer was trained with, and $T_{\mathbf{v}}$ denotes the prompt length used in LTV training. Complete results for different $T_{\mathbf{v}}$ values are provided in Figures 5, 6, and 7 in Appendix D.1.
@@
For regression data, as illustrated in Figure 1, we evaluate the models on classes of linear functions, sparse linear functions, and 2-layer ReLU networks (Garg et al., 2023). While decision trees have also been explored as a function class, they exhibit less sensitivity to positional encodings, which we discuss in Appendix C.2.3.
For NLP tasks, we focus on single-token generation, as considered in previous work (Hendel et al., 2023; Todd et al., 2024). The tasks are derived from widely studied natural language corpora, including SST-2 (Socher et al., 2013a) and CoNLL-2003 (Tjong Kim Sang and De Meulder, 2003). Further details are provided in Appendix C.3.1.
@@
For functional regression, we include FV (Todd et al., 2024) and ICV (Liu et al., 2024) as baselines for comparison. The other methods discussed in Appendix A were primarily designed for language tasks and are not applicable to regression tasks (see Appendix C.2.6 for details). Although FV and ICV were originally proposed for linguistic applications, we adapted them for regression tasks with several optimizations. Additionally, we consider Low-Rank Optimization (LoRA) (Hu et al., 2022) for further discussion. For language tasks, we also include Task Vector (TV) (Hendel et al., 2023) and State Vector (SV) with inner optimization (Li et al., 2024) alongside FV and ICV.
@@
Details of our experimental setup are provided in Appendix C. The full set of results is available in Appendix D, including: (i) LTV trained with various $T_{\mathbf{v}}$, optimized and tested on chain-of-thought (CoT) prompts, and evaluations under distributional shift; (ii) accuracy scores on 15 NLP benchmarks; and (iii) complete ablation results.
@@ 5.1 Evaluation on Regression Tasks
The loss curves for ICL inference are shown in Figure 3. While FV offers some benefits in linear regression, ICV negatively impacts the vanilla model’s performance. Despite extensive optimizations, these results indicate that task encoding formulations, originally crafted for language, do not generalize well to other modalities.
Dataset Prompt Transformer +LoRA +FV +ICV +SV +TV +LTV (ours)
Abstractive
AG News Zero-shot 0.0 $\pm$ 0.0 9.1 $\pm$ 0.2 30.1 $\pm$ 5.7 33.6 $\pm$ 5.6 31.7 $\pm$ 5.7 34.5 $\pm$ 5.0 46.9 $\pm$ 6.2
Few-shot 67.2 $\pm$ 5.8 69.3 $\pm$ 3.8 66.0 $\pm$ 5.8 64.7 $\pm$ 5.4 66.1 $\pm$ 4.5 57.9 $\pm$ 5.8 71.1 $\pm$ 5.6
CommonsenseQA Zero-shot 10.2 $\pm$ 3.7 24.8 $\pm$ 4.1 23.0 $\pm$ 5.2 28.9 $\pm$ 5.2 21.7 $\pm$ 1.1 22.0 $\pm$ 4.0 41.0 $\pm$ 6.1
Few-shot 18.0 $\pm$ 4.7 18.6 $\pm$ 5.5 20.3 $\pm$ 5.0 20.0 $\pm$ 4.6 20.3 $\pm$ 2.1 19.5 $\pm$ 3.4 21.1 $\pm$ 5.0
Sentiment Analysis Zero-shot 0.0 $\pm$ 0.0 0.7 $\pm$ 0.6 0.0 $\pm$ 0.0 4.8 $\pm$ 1.9 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 10.9 $\pm$ 3.8
Few-shot 74.6 $\pm$ 5.4 75.9 $\pm$ 4.1 69.5 $\pm$ 5.7 72.5 $\pm$ 3.2 74.2 $\pm$ 2.5 77.1 $\pm$ 3.4 94.5 $\pm$ 2.8
Extractive
NER-person Zero-shot 5.5 $\pm$ 2.8 15.7 $\pm$ 1.1 48.8 $\pm$ 6.2 49.9 $\pm$ 5.8 46.3 $\pm$ 1.3 47.8 $\pm$ 5.3 58.2 $\pm$ 6.1
Few-shot 11.7 $\pm$ 4.0 12.1 $\pm$ 1.6 56.2 $\pm$ 6.1 65.9 $\pm$ 5.2 53.7 $\pm$ 3.3 62.6 $\pm$ 3.9 79.3 $\pm$ 5.0
NER-location Zero-shot 6.6 $\pm$ 3.1 6.5 $\pm$ 2.4 37.9 $\pm$ 6.0 26.4 $\pm$ 5.1 34.6 $\pm$ 3.9 27.2 $\pm$ 4.7 27.3 $\pm$ 5.5
Few-shot 23.4 $\pm$ 5.2 29.5 $\pm$ 5.5 43.4 $\pm$ 6.1 47.4 $\pm$ 5.9 43.5 $\pm$ 4.4 41.9 $\pm$ 5.6 57.4 $\pm$ 6.1
NER-organization Zero-shot 21.9 $\pm$ 5.1 21.9 $\pm$ 6.6 48.8 $\pm$ 6.2 52.5 $\pm$ 6.0 46.1 $\pm$ 1.2 46.6 $\pm$ 4.3 64.1 $\pm$ 5.9
Few-shot 16.8 $\pm$ 4.6 19.2 $\pm$ 3.8 50.8 $\pm$ 6.2 54.1 $\pm$ 5.0 53.5 $\pm$ 2.6 51.1 $\pm$ 4.1 75.0 $\pm$ 5.3
Table 1: Accuracy scores (%) for zero-shot and few-shot (5-shot) predictions, averaged across 256 random seeds. $\pm$ represents the margin of error at a 95% confidence level. The highest accuracy is marked in boldface, and the statistically best-performing method is highlighted. The complete results are provided in Table 2 in Appendix D.4, along with references to the benchmark datasets.
LTV preserves task information in the OOD case.
LTV shows minimal performance difference when trained with prompt lengths close to $T_{\text{train}}$. However, as $T_{\mathbf{v}}$ increases, LTV’s impact becomes more prominent, achieving optimal performance comparable to the vanilla model at $T=T_{\text{train}}$ (see Appendix D.1 for detailed plots). While this might raise concerns about the validity of the results, as LTV is causally optimized in the OOD cases ($T>T_{\text{train}}$), training LTV with prompt lengths slightly exceeding $T_{\text{train}}$—specifically, $T_{\mathbf{v}}=1.37\times T_{\text{train}}$ for linear functions and $T_{\mathbf{v}}=1.25\times T_{\text{train}}$ for 2-layer ReLU networks—proves sufficient to maintain performance. Extending $T_{\mathbf{v}}$ further offers no significant improvements, indicating that only a small portion of the OOD case is needed for LTV to generalize. This suggests that the attention head weights naturally converge to their optimal values with sufficient training data.
@@
The effects of LTV in regression tasks may resemble those of parameter-efficient fine-tuning (PEFT) methods, as only a small number of parameters ($|\Phi|=L\times J$) are trained to adapt the model to previously unlearned scenarios. To evaluate this, we compare LTV with LoRA (rank 8). While LoRA fine-tunes approximately 196K of the 9.5M GPT-2 parameters, LTV optimizes only 96. Despite LoRA fine-tuning a significantly larger number of parameters on sequences with a maximum length $T_{\text{max}}$, no notable performance improvement is observed. Since LoRA depends on the pre-trained model’s existing positional representations, it cannot address untrained positional encodings. This observation indicates that the task fidelity achieved by LTV is not a result of sole fine-tuning but rather of effective task encoding that can steer the behavior of the model.
@@
No regularization, optimization techniques, dropout, or specific weight initialization were applied to learn the LTV weights. The weights are unbounded and sampled from the standard Gaussian distribution. Visualizing the learned weights as heatmaps revealed they consistently lie within the interval [-3, 3] across tasks. However, no meaningful patterns were observed though, suggesting that any underlying structure may be subtle and challenging to detect visually. Notably, some weights converge to near-zero, supporting the idea that only a subset of attention heads encodes task information. Our method thus provides a principled way to identify and weigh this subset, rather than assuming uniform importance or selecting it manually.
@@ 5.2 Evaluation on Language Tasks
The accuracy scores are reported in Tables 1 and 2. While Todd et al. (2024) primarily considered “filtered” accuracies, which take into account only the test queries where at least one model responds correctly, we present unfiltered accuracies as a fairer metric, counting all samples regardless of model performance.
@@
In 29 out of 30 evaluations (15 tasks with zero- and few-shot prompts), LTV achieves the highest accuracy and ranks among the statistically best-performing methods (see Table 2 in Appendix D.4 for complete results). Furthermore, in 19 out of 30 evaluations, LTV’s performance is statistically superior to all other techniques. We also observe a notable performance by ICV, particularly when the vanilla transformer performs poorly in the zero-shot setting. This effectiveness is attributed to ICV’s approach of crafting unique task vectors for each layer, similar to our method. Lastly, while evaluating LTV with higher shot counts may reveal diminishing gains, we expect the gains to remain nonzero and notable, as the vanilla model may not achieve further contextualization beyond a saturation point. Overall, these findings confirm that LTV excels in both regression and challenging linguistic tasks, supporting our hypothesis about how ICL tasks are formulated within the transformer architecture.
@@ 5.3 Ablation Studies
The KL divergence scores are reported in Table 3 in Appendix D.5. From the baselines, we only include FV, as it is the only method that does not degrade the vanilla transformer. The corresponding probability densities of the last hidden state distributions, visualized through histograms, are also included in Appendix D.5. Additionally, a detailed description of our experimental methods, along with a figure illustrating the pipeline, is provided in Appendix C.6.
@@
We observe that the divergence between the last hidden state distributions of the LTV-integrated model and the optimal-performing model is generally low and significantly decreases when the LTV training length matches or exceeds the middle length. This suggests that LTV’s effectiveness primarily stems from its ability to dominate the last hidden state and align it with that of the optimal-performing model, thereby facilitating task steering. This observation further implies that task steering may generally be achieved through this mechanism, not exclusively by our method. However, a deeper analysis is required to fully understand this process, which is beyond the scope of this study.